# Zero-Shot Floor-Aware Object-Goal Navigation via LLM-Driven Coarse-to-Fine Exploration

### v1

@@ C.2 Implementation Details of Multi-Floor Topology Modeling
Downward stair recognition poses additional challenges. Since stairs are not directly visible from a horizontal viewpoint ($\phi=0^{\circ}$), many existing methods fail to detect them accurately. As a result, they often produce incorrect floor-level identification and cause semantic map mixing. To address this issue, we employ depth anomaly analysis. Projecting depth measurements onto a ground plane reveals regions where points fall below the estimated floor level and exhibit abrupt depth discontinuities. These anomalies trigger active sensing: the agent tilts its camera to $\phi=-30^{\circ}$ and approaches for direct observation. The same cross-modal validation used for ascending stairs confirms descending transitions. This leads to retrieval or creation of lower-floor maps in $\mathcal{L}_{\mathrm{obs}}$, ensuring bidirectional topological consistency.
@@
To ensure safe traversal after reaching stair frontier, we introduce a dynamic waypoint positioned at a distance ahead of the robot. This waypoint is continuously updated based on the robot’s pose to guide the agent through narrow or occluded regions where depth information becomes unreliable. Once the entire robot exits the semantic stair region, it resets its pitch angle and initializes a new floor map if necessary.
@@
Moreover, the BEV maps for unexplored floors are initialized only upon verified stair transitions, with dedicated handling for ascending ($\mathcal{M}^{(n+1)}$) or descending ($\mathcal{M}^{(n-1)}$) cases. This strategy avoids redundant mapping and significantly reduces memory overhead by eliminating the need for full 3D reconstruction. All maps are aligned within a global polar coordinate system $(\rho,\theta,z)$, therefore this unified workflow ensures seamless 3D environment awareness while strictly adhering to RGB-D and 2D odometry constraints.
@@ C.4 Cases Analysis for Stair Ascending
Figure 9: Cases Analysis for Stair Ascending. Top row: After traversaling the current floor, the agent makes a multi-floor decision to navigate upstairs and successfully find the goal on a higher floor in the HM3D dataset. Bottom row: The agent actively infers that the target is on a higher floor and decides to ascend, successfully navigating to the goal on a higher floor in the MP3D dataset.
@@
Initially, the agent operates under the assumption of a single floor, using BEV map lists $\mathcal{L}_{obs}=[\mathcal{M}_{obs}^{(1)}]$ and $\mathcal{L}_{val}=[\mathcal{M}_{val}^{(1)}]$, performing intra-floor navigation until detecting an up-stair region through both staircase detection and semantic segmentation. Upon detection, it initializes maps for the higher floor ($\mathcal{M}_{obs}^{(2)}$, $\mathcal{M}_{val}^{(2)}$).
The agent initiates upstairs transition when either: (1) the current floor exploration completes while higher unexplored floors remain, as shown in the example of the top row in Fig. 9; or (2) the step count exceeds $\tau_{floor}$ steps and the time since the last floor decision exceeds $\tau_{period}$ steps (triggering the LLM’s multi-floor decision), as shown in the example of the bottom row in Fig. 9.
@@ C.5 Cases Analysis for Stair Descending
Figure 10: Cases Analysis for Stair Descending. Top row: After traversaling the current floor, the agent makes a multi-floor decision to navigate downstairs and successfully find the goal on a lower floor in the HM3D dataset. Bottom row: The agent actively infers that the target is on a lower floor and decides to descend, successfully navigating to the goal on a lower floor in the MP3D dataset.
@@
The agent initiates downstairs transition when either: (1) the current floor exploration completes while lower unexplored floors remain, as shown in the example of the top row in Fig. 10; or (2) the step count exceeds $\tau_{floor}$ steps and the time since the last floor decision exceeds $\tau_{period}$ steps (triggering the LLM’s multi-floor decision), as shown in the example of the bottom row in Fig. 10.
@@
After descending the stairs, the agent initializes the navigation stage on the new floor while establishing floor connectivity through bidirectional stair mapping using the BEV maps $\mathcal{M}_{obs}^{(1)}$ and $\mathcal{M}_{val}^{(1)}$. The lower floor’s up-stair parameters inherit the higher floor’s down-stair geometry with start/end points inverted while preserving the centroid. Navigation then continues with intra-floor rules, periodically invoking the LLM decision when unexplored floors remain in the exploration state.
@@ C.6 Cases Analysis for Stairwells
The agent may initialize its navigation within a stairwell, either positioned on a landing between stair flights (Fig. 11, top) or directly inside the stair section (Fig. 11, bottom). Due to limited initial contextal information of scenario environment, we incorporate prior floor knowledge with the specific goal into the LLM. When the agent detects both ascending and descending staircases, this prior knowledge informs the LLM’s multi-floor decision-making process, enabling probabilistic estimation of the target’s vertical location (upstairs vs. downstairs). This optimization significantly improves navigation efficiency.
@@ C.7 Cases Analysis for Floor Revisiting
As shown in the Fig. 12, these cases demonstrate the effectiveness and robustness of our algorithm in multi-floor navigation.
The top row of examples is from the scene of the HM3D dataset. The agent initially leverages prior floor knowledge to make multi-floor decisions. After an unsuccessful search on the higher floor, the system actively reevaluates its strategy and successfully relocates the target might on the lower floor through timely floor revisit. This adaptive behavior demonstrates the algorithm’s robustness in recovering from initial decision inaccuracies.
The bottom row presents a more complex case with dual staircases connecting the first and second floors from the MP3D dataset. The scene has two staircases between the 1st floor and the 2nd floor. Following partial exploration of the first floor, the agent prioritizes searching the second floor based on prior probability estimates. After searching for a while on the 2nd floor but does not find the target object, the agent needs to choose to go up to the 3rd floor or down to the 1st floor. Because the prior knowledge suggests that the probability of finding the target on the 1st floor is higher than that on the 3rd floor, the agent revisits the 1st floor from another staircase and successfully reaches the target, which validates both the effectiveness of our probability-driven approach and the algorithm’s capability to handle architectural complexities.
Appendix D Limitations
Table 8: Stair detection performance in a subset of HM3D multi-floor scenes
@@ Appendix D Limitations
While our method demonstrates effective multi-floor navigation capabilities, several limitations merit discussion. First, the accuracy of stair detection and floor perception remains imperfect, as evidenced shown in Tab. 8 by our evaluation on HM3D multi-floor episodes. The current system achieves 66.7% correct stair detection, but suffers from 20% false positives where non-stair structures are misclassified as stairs or stair directions are inverted, along with 13.3% missed detections where existing stairs go unrecognized. These errors primarily emerge from challenging visual conditions such as occluded stairwells, ambiguous depth signatures in corner cases, and visual similarities between stairs and other architectural features like ramps or platforms. Although our floor revisiting mechanism provides robustness against such detection failures, future improvements could explore multi-modal sensor fusion or temporal consistency checks to enhance reliability.
@@
Appendix E Prompt Design
@@
These components are meticulously designed to guide the model in leveraging both probabilistic data and environmental context for zero-shot multi-floor object navigation. The structured format not only ensures clarity and consistency in the model’s responses but also facilitates effective communication and decision-making in unseen environments. By providing a comprehensive framework, the prompts enable the model to perform complex reasoning tasks with precision and reliability, which is critical for zero-shot navigation applications requiring high accuracy and interpretability.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Both contributions and scope have been discussed in abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The detailed analysis on limitations have been discussed in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This is an empirical study that excludes theory assumptions and proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All information needed to reproduce the experimental results have been disclosed. To ensure reproducibility, code and data are committed to be publicly available.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The detailed implementation procedures have been included in the appendix. To ensure reproducibility, code and data are committed to be publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All training and test details have been discussed in either main body or appendix. To ensure reproducibility, code and data are committed to be publicly available.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Sufficient information about experiment settings have been discussed.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The details on computing resources have been discussed.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This research follows the NeurIPS Code of Ethics properly.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The discussion on societal impacts has been included in the appendix.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: The discussion on safeguards has been included in the appendix.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The acknowledgments on licenses have been included in the appendix.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The discussions on new assets have been included in the appendix.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ C.2 Implementation Details of Multi-Floor Topology Modeling
Downward stair recognition poses additional challenges. Since stairs are not directly visible from a horizontal viewpoint ($\phi=0^{\circ}$), many existing methods fail to detect them accurately. As a result, they often produce incorrect floor-level identification and cause semantic map mixing. To address this issue, we employ depth anomaly analysis. Projecting depth measurements onto a ground plane reveals regions where points fall below the estimated floor level and exhibit abrupt depth discontinuities. These anomalies trigger active sensing: the agent tilts its camera to $\phi=-30^{\circ}$ and approaches for direct observation. The same cross-modal validation used for ascending stairs confirms descending transitions. This leads to the retrieval or creation of lower-floor maps in $\mathcal{L}_{\mathrm{obs}}$, ensuring bidirectional topological consistency.
@@
To ensure safe traversal after reaching the stair frontier, we introduce a dynamic waypoint positioned at a distance ahead of the robot. This waypoint is continuously updated based on the robot’s pose to guide the agent through narrow or occluded regions where depth information becomes unreliable. Once the entire robot exits the semantic stair region, it resets its pitch angle and initializes a new floor map if necessary.
@@
Moreover, the BEV maps for unexplored floors are initialized only upon verified stair transitions, with dedicated handling for ascending ($\mathcal{M}^{(n+1)}$) or descending ($\mathcal{M}^{(n-1)}$) cases. This strategy avoids redundant mapping and significantly reduces memory overhead by eliminating the need for full 3D reconstruction. All maps are aligned within a global polar coordinate system $(\rho,\theta,z)$, therefore, this unified workflow ensures seamless 3D environment awareness while strictly adhering to RGB-D and 2D odometry constraints.
@@ C.3 Stair Detection Visualization for Multi-Floor Navigation
Figure 9: Cases Analysis for Stair Ascending. Top row: After traversing the current floor, the agent makes a multi-floor decision to navigate upstairs and successfully find the goal on a higher floor in the HM3D dataset. Bottom row: The agent actively infers that the target is on a higher floor and decides to ascend, successfully navigating to the goal on a higher floor in the MP3D dataset.
@@ C.4 Cases Analysis for Stair Ascending
Initially, the agent operates under the assumption of a single floor, using BEV map lists $\mathcal{L}_{obs}=[\mathcal{M}_{obs}^{(1)}]$ and $\mathcal{L}_{val}=[\mathcal{M}_{val}^{(1)}]$, performing intra-floor navigation until detecting an upstairs region through both staircase detection and semantic segmentation. Upon detection, it initializes maps for the higher floor ($\mathcal{M}_{obs}^{(2)}$, $\mathcal{M}_{val}^{(2)}$).
The agent initiates upstairs transition when either: (1) the current floor exploration completes while higher unexplored floors remain, as shown in the example of the top row in Fig. 9; or (2) the step count exceeds $\tau_{\mathrm{floor}}$ steps and the time since the last floor decision exceeds $\tau_{\mathrm{period}}$ steps (triggering the LLM’s multi-floor decision), as shown in the example of the bottom row in Fig. 9.
@@
Figure 10: Cases Analysis for Stair Descending. Top row: After traversing the current floor, the agent makes a multi-floor decision to navigate downstairs and successfully find the goal on a lower floor in the HM3D dataset. Bottom row: The agent actively infers that the target is on a lower floor and decides to descend, successfully navigating to the goal on a lower floor in the MP3D dataset.
@@ C.5 Cases Analysis for Stair Descending
The agent initiates downstairs transition when either: (1) the current floor exploration completes while lower unexplored floors remain, as shown in the example of the top row in Fig. 10; or (2) the step count exceeds $\tau_{\mathrm{floor}}$ steps and the time since the last floor decision exceeds $\tau_{\mathrm{period}}$ steps (triggering the LLM’s multi-floor decision), as shown in the example of the bottom row in Fig. 10.
@@
After descending the stairs, the agent initializes the navigation stage on the new floor while establishing floor connectivity through bidirectional stair mapping using the BEV maps $\mathcal{M}_{obs}^{(1)}$ and $\mathcal{M}_{val}^{(1)}$. The lower floor’s upstairs parameters inherit the higher floor’s down-stair geometry with start/end points inverted while preserving the centroid. Navigation then continues with intra-floor rules, periodically invoking the LLM decision when unexplored floors remain in the exploration state.
@@ C.6 Cases Analysis for Stairwells
The agent may initialize its navigation within a stairwell, either positioned on a landing between stair flights (Fig. 11, top) or directly inside the stair section (Fig. 11, bottom). Due to limited initial contextual information of the scenario environment, we incorporate prior floor knowledge with the specific goal into the LLM. When the agent detects both ascending and descending staircases, this prior knowledge informs the LLM’s multi-floor decision-making process, enabling probabilistic estimation of the target’s vertical location (upstairs vs. downstairs). This optimization significantly improves navigation efficiency.
@@ C.7 Cases Analysis for Floor Revisiting
As shown in Fig. 12, these cases demonstrate the effectiveness and robustness of our algorithm in multi-floor navigation.
The top row of examples is from the scene of the HM3D dataset. The agent initially leverages prior floor knowledge to make multi-floor decisions. After an unsuccessful search on the higher floor, the system actively reevaluates its strategy and successfully relocates the target might on the lower floor through a timely floor revisit. This adaptive behavior demonstrates the algorithm’s robustness in recovering from initial decision inaccuracies.
The bottom row presents a more complex case with dual staircases connecting the first and second floors from the MP3D dataset. The scene has two staircases between the 1st floor and the $2$nd floor. Following partial exploration of the first floor, the agent prioritizes searching the second floor based on prior probability estimates. After searching for a while on the $2$nd floor but not finding the target object, the agent needs to choose to go up to the $3$rd floor or down to the 1st floor. Because the prior knowledge suggests that the probability of finding the target on the 1st floor is higher than that on the 3rd floor, the agent revisits the $1$st floor from another staircase and successfully reaches the target, which validates both the effectiveness of our probability-driven approach and the algorithm’s capability to handle architectural complexities.
Appendix D Discussions & Limitations
Table 8: Stair detection performance in a subset of the HM3D multi-floor scenes.
@@ Appendix D Discussions & Limitations
While our method demonstrates effective multi-floor navigation capabilities, several limitations merit discussion. First, the accuracy of stair detection and floor perception remains imperfect, as evidenced shown in Tab. 8 by our evaluation on HM3D multi-floor episodes. The current system achieves $66.7\%$ correct stair detection, but suffers from $20\%$ false positives where non-stair structures are misclassified as stairs or stair directions are inverted, along with $13.3\%$ missed detections where existing stairs go unrecognized. These errors primarily emerge from challenging visual conditions such as occluded stairwells, ambiguous depth signatures in corner cases, and visual similarities between stairs and other architectural features like ramps or platforms. Although our floor revisiting mechanism provides robustness against such detection failures, future improvements could explore multi-modal sensor fusion or temporal consistency checks to enhance reliability.
@@
Appendix E Detailed Prompt Design
@@
These components are meticulously designed to guide the model in leveraging both probabilistic data and environmental context for zero-shot multi-floor object navigation. The structured format not only ensures clarity and consistency in the model responses but also facilitates effective communication and decision-making in unseen environments. By providing a comprehensive framework, the prompts enable the model to perform complex reasoning tasks with precision and reliability, which is critical for zero-shot navigation applications requiring high accuracy and interpretability.